# Generative Emergent Communication: Large Language Model is a Collective World Model

### v1

@@ II-B Generative EmCom
II-C Language Game as Decentralized Bayesian Inference
@@
    Perception: Speaker and listener agents (Sp and Li) observe the $d$-th object, obtain $x^{Sp}_{d}$, and $x^{Li}_{d}$ infers their internal representations $z_{d}^{Sp}$ and $z_{d}^{Li}$, respectively.
@@
It has been demonstrated that the MHNG is equivalent to the Metropolis-Hastings algorithm for inferring latent variables in a probabilistic generative model. This model conditions the internal representations $z^{k}$ of multiple agents, acting as representation learning machines, on a common external representation $m$. Although the original study assumed two agents and a categorical message $m$, the core PGM of generative EmCom underlying this theory does not make these assumptions. Consequently, this fundamental idea can be extended in various ways.
As described, unlike referential games, MHNG assumes a joint attention performed by two agents. This assumption may seem strange from the game-theoretic approach to EmCom, such as referential games. However, developmental studies suggest that joint attention is fundamental to human infant language acquisition. Human infants develop joint attention capabilities before vocabulary explosion occurs, and joint attention serves as a crucial foundation for language acquisition [7]. Gergely et al. highlighted that when children incorporate parental instructions into their learning process, they presuppose that parents have the intention to teach them [53].
Moreover, interactions modeled by referential games—where agents identify objects and receive feedback—are rarely observed in natural language acquisition in cognitive development. Unlike game-theoretic approaches, MHNG, as an example of generative EmCom, assumes cooperative behavior in symbol emergence, which aligns with scholarly observations about human language acquisition [7]. Thus, although conventional (discriminative) EmCom provides rational models for both cooperative and competitive games, the assumptions of generative EmCom better align with empirical findings in human communication behavior.
@@
The original idea of MH-based EmCom was introduced by Hagiwara et al. [48] and later generalized and formalized by Taniguchi et al. [47], who clarified its connection to representation learning. The concept of MHNG has since been extended and validated in various ways. Furukawa et al. [54] demonstrated that similar symbol emergence can be observed even when the relationship between $m$ and $z$ is head-to-head instead of tail-to-tail. The extension to multimodal sensory information was achieved by Hagiwara et al. [55], who demonstrated that MHNG can lead to symbol emergence even when agents have different sensory modalities. Their work also showed that modality information possessed by other agents, but not by oneself, can facilitate object category formation. Similar multimodal extensions have been explored with variational autoencoder (VAE)-based representation learning by Hoang et al. [56], though their work suggests that the mechanism for integrating multimodal sensory information influences symbol emergence.
Although these MHNG studies involve two agents, a mathematical extension to N-agent conditions was developed by Inukai et al. [57], who introduced a recursive structure in communication while maintaining its characteristics as distributed Bayesian inference. They argued that random partner selection for MHNG can be considered as a one-sample and limited-length approximation of this approach.
Although several generative EmCom studies define the sign $m$ as a categorical variable, this is not a necessary assumption of the framework. A recent study by Hoang et al. [58] demonstrates that sharing compositional word sequences is possible within the MHNG framework, similar to numerous EmCom studies. From Peirce’s semiotic perspective, signs include both compositional discrete sequences and continuous signs such as voice pitch and facial expressions. A recent study by You et al. [59] demonstrated that such continuous signs can emerge within the generative EmCom framework, whereas Saito et al. [60] modeled the emergence of compositional signs from continuous time-series information as signals.
@@
MHNG serves as a basic language game to represent symbol emergence by realizing distributed Bayesian inference and inferring latent variables of the generative model, corresponding to language. The critical point is that the inference of the posterior distribution is performed in a decentralized manner through language games. This suggests that being based on the MH method is not a necessary condition. The MHNG-based approach represents an initial step in modeling symbol emergence (or language emergence, EmCom) as decentralized Bayesian inference. Future studies should explore the possibility of constructing generative EmCom models by distributing various inference methods (e.g., [62]).
III EmCom for Multi-agent Cooperation
III-A Generative EmCom for Multi-agent Reinforcement Learning
@@
In contrast, generative EmCom allows us to formulate MARL with emergent communication in a Bayesian manner by incorporating the idea of CaI [76], which is a theory to formulate RL as a PGM [77, 78]. Figure 3 shows a graphical model of cooperative action generation between two agents, and the details of each stochastic variable are listed in Table I. The behavior of each agent is generated through a Markov decision process with a prior variable $m_{t}$. The state $z^{k}_{t}$ of an agent at time $t$ is determined according to state $z^{k}_{t-1}$, action $a^{k}_{t-1}$, and message $m_{t}$, which is the shared latent variable:
@@
For the reward of the group of agents $r^{AB}$, the following approximate reward models are learned by each agent:
$p(o^{AB}_{t}=1|z^{k}_{t},a^{k}_{t})\approx\propto\exp(r^{AB}(z^{A}_{t},a^{A}_{t},z^{B}_{t},a^{B}_{t})).$ (5)
[…]
@@
Equation (7) describes state planning, which can be computed based on the CaI framework [76]. Equation (8) describes the inference of the message and can be formulated using the MHNG proposed by Taniguchi et al. [55, 47], which allows both agents to infer messages through communication without observing each other’s internal states. Note that MHNG enables two agents to perform sampling in (8) without simultaneous observations of $z^{A}_{t},z^{B}_{t}$.
Thus, MHNG can be used not only for multimodal object categorization and naming but also for action coordination among multi-agents using PGM to formulate MARL, that is, modeling EmCom for multi-agent cooperation.
III-B Generative EmCom on World models
The generative EmCom for MARL can be extended to involve representation learning in the same way as we discussed in Section 2.3. This extension implies the integration of the idea of world models into generative EmCom.
The concept of a world model represents an internal model within an agent that captures the dynamics of environmental states, their responses to the actions of the agent, and their relationships with sensory inputs [79, 18, 19]. The concept of world model has its origins in the early days of AI and robotics studies [80]. Initial studies on ML investigated techniques for agents to autonomously construct and adapt their world models [81, 82]. Currently, the term generally refers to predictive frameworks [83], predominantly implemented using DNN architectures.
The concept of world models is closely related to the idea of PCs. PC is the idea that the brain constantly predicts sensory information and updates its internal models to minimize prediction errors. In the context of cognitive robotics and AI, world models provide the structure for representing and reasoning about the environment, whereas PC offers a mechanism for learning and updating these models based on sensory experiences. The free energy principle and active inference further unify these concepts, suggesting that both perception and action can be considered as processes of minimizing prediction errors or free energy. A theoretical connection exists between them[19] .
State transition: $\quad z_{t+1}\sim p(z_{t+1}|z_{t},a_{t})$ (9)
Observation: $\quad x_{t}\sim p(x_{t}|z_{t})$ (10)
Inference: $\quad z_{t}\sim q(z_{t}|x_{1:t},a_{1:t-1})$ (11)
Incorporating world models into MARL based on generative EmCom is natural and reasonable. By estimating $z_{t}$ using the inference model of a world model $q(z_{t}|x_{1:t},a_{1:t-1})$, an agent can estimate its state. Agents are expected to perform EmCom in the same manner as described in Section 3.1. Figure 4 shows a PGM of generative EmCom on POMDPs. This demonstrates the theoretical connection between EmCom or language and world models. When considered in this way, CPC is the theory explaining that language is something that integrates world models of populated agents. This will be discussed in more detail in Section 5.
@@ III-B Generative EmCom on World models
Several studies linking EmCom and world models have been conducted. Recent studies on EmLang have focused on MARL approaches, aiming to autonomously develop languages to solve tasks [2]. Because each agent typically has access only to its own observations, inferring the underlying state of the environment becomes crucial. Consequently, researchers have explored enabling RL agents to develop EmLangs for inter-agent communication while simultaneously learning world models to infer environmental states. Notably, not all MARL communication studies focus on developing symbol-based languages; some treat communication as simple information exchange. This distinction separates “ EmLang” from “learning tasks with communication.” [49]
Alexander et al. proposed a message-conditioned world model that improves the quality of communication among agents, facilitating the emergence of meaningful messages and enhancing the interpretability of the model [84]. Similarly, Kenzo Lobos-Tsunekawa et al. extended the widely used Dreamer world model by introducing a global world model, used exclusively during training. By incorporating information from all agents, they achieved improved sample efficiency through shared imagination [85]. Toledo et al. also used Dreamer but introduced GNNs to facilitate communication; however, their work is classified under ”learning tasks with communication” [86] because it does not aim to develop a model of EmLang.
IV Relationship with Conventional EmCom
Recent years have witnessed an increasing presence of research papers focused on EmCom at ML conferences such as NeurIPS and ICLR. This trend underscores the renewed interest in EmCom. In the field of EmCom, Lewis’ signaling game [87] framework is often used as a model of communication while other various formulations have also been proposed [35, 36, 37]. The game players, or agents, are optimized to achieve some goals, resulting in the communication protocols, which are referred to as EmLang. In the EmCom field, the extent to which EmLangs resemble human languages is frequently questioned from the perspectives of evolutionary and/or computational linguistics. For instance, questions have been raised about whether EmLangs have compositionality [88, 89] or whether they follow the well-known statistical universal properties of natural languages [90, 91, 92, 93]. Additionally, cases exist where cognitive constraints were modeled to induce characteristics of natural languages [94, 95].

### v2

@@ 3.2 Formalizing Generative EmCom as Decentralized Bayesian Inference
3.3 The Naming Game as a Microcosm of Collective Inference
@@
    Perception: Speaker and listener agents (Sp and Li) observe the $d$-th object, obtain $x^{Sp}_{d}$, and $x^{Li}_{d}$, and infer their internal representations $z_{d}^{Sp}$ and $z_{d}^{Li}$, respectively.
@@
It has been demonstrated that the MHNG is equivalent to the Metropolis-Hastings algorithm for inferring latent variables in a probabilistic generative model. As a result, the MHNG protocol guarantees that the CFE, defined by the KL-divergence between the agents’ collective belief and the true posterior, is monotonically non-increasing in expectation. This is because the underlying distributed Metropolis-Hastings sampler satisfies the detailed balance condition, a well-established result related to the Data Processing Inequality [21], ensuring convergence towards the target posterior distribution. This model conditions the internal representations $z^{k}$ of multiple agents, acting as representation learning machines, on a common external representation $m$. Although the original study assumed two agents and a categorical message $m$, the core probabilistic graphical model (PGM) of generative EmCom underlying this theory does not make these assumptions. Consequently, this fundamental idea can be extended in various ways.
As described, unlike referential games, MHNG assumes a joint attention performed by two agents [^1] The assumption of joint attention is specific to the naming game, as its primary task is to establish a shared vocabulary for a common referent. However, the broader CPC framework does not depend on this. On the contrary, when applied to general multi-agent cooperation, its core strength lies in integrating different, partially observable perspectives, as we will discuss in Section 4.. This assumption may seem strange from the game-theoretic approach to EmCom, but it is a deliberate choice made to depart from the feedback-dependent principle of success or failure that underpins many such models. Developmental studies suggest that such explicit feedback interactions are rarely observed in natural language acquisition and are an implausible metaphor for how infants learn [126]. For human infants, joint attention is not a consequence of a communicative act, but rather a cognitive foundation that is in place before the vocabulary explosion occurs. Furthermore, children engaging in learning presuppose a pedagogical intention from adults [23], rather than simply testing a communication channel. Therefore, by assuming cooperative behavior grounded in joint attention, MHNG, as an example of generative EmCom, better aligns with empirical findings in human communication behavior than conventional discriminative models.
@@
The original idea of MH-based EmCom was introduced by Hagiwara et al. [44] and later generalized and formalized by Taniguchi et al. [124], who clarified its connection to representation learning. The concept of MHNG has since been extended and validated in various ways. The extension to multimodal sensory information was achieved by Hagiwara et al. [43], who demonstrated that MHNG can lead to symbol emergence even when agents have different sensory modalities. Their work also showed that modality information possessed by other agents, but not by oneself, can facilitate object category formation. Similar multimodal extensions have been explored with variational autoencoder (VAE)-based representation learning by Hoang et al. [48], though their work suggests that the mechanism for integrating multimodal sensory information influences symbol emergence. Although these MHNG studies involve two agents, a mathematical extension to $N$-agent conditions was developed by Inukai et al. [52], who introduced a recursive structure in communication while maintaining its characteristics as distributed Bayesian inference. They argued that random partner selection for MHNG can be considered as a one-sample and limited-length approximation of this approach.
The Generative EmCom framework is general, and its core mechanisms are not limited to the exchange of simple categorical signs; they can be readily extended to handle complex, compositional languages. A recent study by Le Hoang et al. [69] demonstrates that sharing compositional word sequences is possible within the MHNG framework, similar to numerous EmCom studies. Furthermore, Matsui et al. [80] extended this concept to a full captioning game, where multiple Vision-Language Models interactively generate and refine natural language captions for images. From Peirce’s semiotic perspective, signs include both compositional discrete sequences and continuous signs such as voice pitch and facial expressions. A recent study by You et al. [143] demonstrated that such continuous signs can emerge within the generative EmCom framework, whereas Saito et al. [107] modeled the emergence of compositional signs from continuous time-series information as signals.
@@
MHNG serves as a basic language game to represent symbol emergence by realizing distributed Bayesian inference and inferring latent variables of the generative model, corresponding to language. The critical point is that the inference of the posterior distribution is performed in a decentralized manner through language games. This suggests that being based on the MH method is not a necessary condition. The MHNG-based approach represents an initial step in modeling symbol emergence (or language emergence, EmCom) as decentralized Bayesian inference. Future studies should explore the possibility of constructing generative EmCom models by distributing various inference methods (e.g., Hoang et al. [49]).
4 Emergence of a Collective World Model through Multi-Agent Interaction and Cooperation
The previous section used the MHNG to illustrate decentralized inference under the simplified assumption of joint attention. This section expands upon that principle, turning to more general multi-agent systems where joint attention is not required. Here, the essence of the CPC framework becomes clear: to integrate the diverse, partially observable perspectives of multiple agents through communication. We will explore how an emergent language allows agents to resolve their individual uncertainties and coordinate their actions by inferring a shared understanding of the world. This exploration begins with the application of Generative EmCom to MARL, considering that Type 1 world models are theoretically rooted in reinforcement learning studies.
4.1 From Communication to Cooperation in Multi-Agent Systems
@@
In contrast, generative EmCom allows us to formulate MARL with EmCom in a Bayesian manner by incorporating the idea of CaI [71], which is a theory to formulate RL as a PGM [28, 85]. Figure 3 shows a graphical model of cooperative action generation between two agents, and the details of each stochastic variable are listed in Table 1. The behavior of each agent is generated through a Markov decision process with a prior variable $m_{t}$. The state $z^{k}_{t}$ of an agent at time $t$ is determined according to state $z^{k}_{t-1}$, action $a^{k}_{t-1}$, and message $m_{t}$, which is the shared latent variable:
@@
For the group of agents, the true optimality $o^{AB}_{t}$ depends on the joint state and actions, with its probability being proportional to a global reward function $r^{AB}$:
$p(o^{AB}_{t}=1|z^{A}_{t},a^{A}_{t},z^{B}_{t},a^{B}_{t})\propto\exp(r^{AB}(z^{A}_{t},a^{A}_{t},z^{B}_{t},a^{B}_{t})).$ (13)
However, since each agent $k$ cannot observe others’ internal variables, it must learn an approximate model of this group optimality conditioned only on its own state $z^{k}_{t}$ and action $a^{k}_{t}$. Agent $k$ therefore learns its own predictive model of group success as:
$p(o^{AB}_{t}=1|z^{k}_{t},a^{k}_{t})\propto\exp(\hat{r}^{AB}_{k}(z^{k}_{t},a^{k}_{t})),$ (14)
where $\hat{r}^{AB}_{k}(z^{k}_{t},a^{k}_{t})$ is a function learned by agent $k$ that locally approximates the global reward. This formulation eliminates the necessity for an agent to access others’ internal variables to model group success, while still allowing it to contribute to the collective goal.
@@
Equation (16) describes state planning, which can be computed based on the CaI framework [71]. Equation (17) describes the inference of the message and can be formulated using the MHNG proposed by Hagiwara et al. [43] and Taniguchi et al. [124], which allows both agents to infer messages through communication without observing each other’s internal states. Note that MHNG enables two agents to perform sampling in (17) without simultaneous observations of $z^{A}_{t},z^{B}_{t}$.
Thus, MHNG can be used not only for multimodal object categorization and naming but also for action coordination among multi-agents using PGM to formulate MARL, that is, modeling EmCom for multi-agent cooperation. This theoretical connection has been recently instantiated in the MARL-CPC framework, which successfully applies the principles of CPC to on-policy RL algorithms and demonstrates effective communication in non-cooperative settings [141].
@@
4.2 Language as an External Representation for Integrating World Models
As established in Section 2.2, an individual agent’s world model is an internal mechanism used to predict future latent states based on current states and actions, i.e., it learns the state transition dynamics $p(z_{t+1}|z_{t},a_{t})$. In a multi-agent setting where each agent has only partial observations of the environment, each agent learns its own subjective world model. Without a mechanism for exchanging information, these individual world models remain isolated, and the agents lack a way to form a shared, consistent understanding of the global environment.
Our Generative EmCom framework introduces such a mechanism by incorporating a shared latent variable, the message $m$, into the world model of each agent. This fundamentally alters the state transition dynamics. As depicted in the PGM for a single agent with communication (Figure 4), the world model of agent $k$, parameterized by $\theta^{k}$, now learns the dynamics conditioned on the message: $p(z^{k}_{t+1}|z^{k}_{t},a^{k}_{t},m_{t})$. Here, the message $m_{t}$ is not part of the action space; instead, it acts as a conditioning variable that directly influences the agent’s prediction of its future internal state $z^{k}_{t+1}$. This latent state, in turn, informs the agent’s policy, $\pi^{k}$, which generates the subsequent action.
This architecture creates a powerful feedback loop for knowledge integration. As formalized in our framework (Section 3.2), the message $m_{t}$ is inferred from the collective internal states of all agents, $\{z^{k}_{t}\}_{k}$. In turn, this collectively-informed message influences the future internal state of each individual agent. Therefore, the process of learning to use the language $m$ effectively becomes a process of learning to align these individual world models. Each agent learns to encode its own unique knowledge into the shared language and to decode the language to update its own world model with information it could not observe directly. This perspective aligns with recent works that have demonstrated the effectiveness of message-conditioned world models for improving coordination [22, 77, 88].
Through this process, the emergent language $m$ becomes more than just a set of signals for coordinating immediate actions. It becomes the shared, external representation that embodies the integrated knowledge of the individual world models. It is the medium through which a collective of agents, each with a subjective Type 1 World Model, can construct a shared, objective Type 2 World Model (i.e., Model of the World), allowing the group to model the world from a unified perspective that transcends any single agent’s limited view.
4.3 Collective World Model as an Abstractive Prior over Embodied Experiences
To understand how language integrates individual world models into a collective one, we first consider the hierarchy of experiences that language can represent, as illustrated in Figure 5. This hierarchy progresses through three levels of increasing complexity. Level A represents the bidirectional relationship between a static observation, such as an image $x_{k}$, and a linguistic description $m$. This includes both image captioning (inferring $m$ from $x_{k}$) and image generation (generating $x_{k}$ from $m$[134, 139, 102]. Level B extends this to dynamic sequences, where the language $m$ corresponds to a stream of observations over time, $\{x^{k}_{t}\}_{t=1:T}$, as seen in video captioning and generation tasks [132, 140]. The hierarchy culminates in Level C, which models the full sensorimotor loop of an embodied agent. Here, language $m$ is grounded in both dynamic perception $\{x^{k}_{t}\}_{t=1:T}$ and a sequence of actions $\{a^{k}_{t}\}_{t=1:T}$. This final level corresponds to the world models used in modern robotics, often called Vision-Language-Action (VLA) models, which learn from an agent’s own embodied interactions [5, 61, 25, 145, 58].
Crucially, in a collective setting, the shared message $m$ is inferred from the observations of all $K$ agents, e.g., $m\sim q(m|\{x^{k}_{1:T}\}_{k})$. This collective inference process itself constitutes a form of (external) representation learning, where the structure of the shared language $m$ comes to reflect the integrated knowledge of the entire group.
The generative process for each of these three levels can be explicitly formulated as a Probabilistic Graphical Model (PGM), as shown in Figure 6. At Level A, the PGM captures the relationship between a static image observation $x^{k}$ and a message $m$, mediated by an internal representation $z^{k}$. At Level B, this model is extended to handle temporal dynamics, where a sequence of video frames $\{x^{k}_{1:T}\}$ informs a sequence of internal states $\{z^{k}_{1:T}\}$. As shown in the PGMs for Level A and Level B (Figure 6 (A) and (B)), the processes are governed by two sets of parameters for each agent $k$: $\theta^{k}$ represents the parameters of the agent’s internal model (i.e., the VLM), while $\phi^{k}$ denotes the parameters of the prior over the shared message $m$. Finally, Level C depicts the complete generative model for an embodied agent, where both sensory observations $\{x^{k}_{1:T}\}$ and actions $\{a^{k}_{1:T}\}$ jointly determine the agent’s internal state trajectory. The parameters of this comprehensive model, $\bar{\theta}^{k}$, therefore encompass not only the agent’s predictive world model but also its policy. These PGMs represent concrete instances of the overarching CPC hypothesis.
This principle generalizes directly as we move up the hierarchy to a collective of fully embodied VLA agents (Level C). Here, the framework models a population of agents, each possessing its own sophisticated world model for navigating its environment (Figure 6 (C)). The interactive, decentralized process of these agents developing a shared language to communicate about their sensorimotor experiences ($(\{x^{k}_{1:T}\}_{k},\{a^{k}_{1:T}\}_{k})$ becomes mathematically equivalent to the process of learning a single, abstract, hierarchical world model that spans the entire group. Therefore, we argue that embodied symbol emergence, under the CPC framework, is collective world modeling.
The emergent language $m$ is not merely a tool for this process; it is the very instantiation of the Collective World Model. It functions as a highly abstract structured prior over the entire space of possible embodied experiences within the society. The dynamic, interactive process of collective sense-making described above does not simply vanish; it leaves behind a static artifact in the form of a text corpus. We can think of this corpus as a “fossil record” of a society’s ongoing, collective inference about the world [^2] This idea has also been applied to model the process of scientific activity itself from the perspective of CPC, a framework known as CPC as a model of science (CPC-MS) [122].. From this perspective, each sentence or document within a corpus can be viewed as a sample from an underlying posterior distribution. More formally, we consider a sentence $m_{j}^{[i]}$ describing a particular event or situation $j$ to be a sample from the approximate posterior $q(m|\{\{x^{k}_{t}\}_{t=1:T},\{a^{k}_{t}\}_{t=1:T}\}_{k})$. It is this product, rich with the implicitly encoded structure of the collective world model, that serves as the training data for LLMs.
5 LLM as Collective World models
5.1 Language Corpora as Samples from the Collective World Model
The preceding sections have established our central premise: that human language, as a product of CPC, can be understood as a Collective World Model externalized in a shared symbolic system. Building on this foundation, we now connect this theory to the training of LLMs.
As argued in Section 4.3, each sentence $m$ within a text corpus is not an isolated artifact but can be viewed as a sample drawn from a complex posterior distribution, conditioned on the unobserved, collective sensorimotor experiences of the society that generated it:
$m\sim q(m|\{\{x^{k}_{t}\}_{t=1:T},\{a^{k}_{t}\}_{t=1:T}\}_{k})$ (18)
The vast web-scale corpora used to train modern LLMs, therefore, represent a massive dataset of samples drawn from this collective, world-grounded inference process. The fundamental objective of an LLM during pre-training, typically next-token prediction, is an autoregressive method to learn a model, which we denote as $p_{\text{LM}}(m)$, that approximates the marginal distribution $p(m)$ of these textual samples. From this perspective, the task of an LLM is not merely to learn statistical correlations in text, but to implicitly model the output of the complex, embodied, and collective process that produced the text in the first place.
5.2 The LLM as a Reconstructor of Collective Representation
As established in the previous section, LLMs are trained to model the distribution $P(m)$ of language corpora, which are themselves samples from a collective, world-grounded process. When a trained LLM processes a given sentence $m$, we can posit that it forms a corresponding high-level internal representation or latent state, which we denote as $z_{\text{LM}}$. This constitutes an inference or recognition process within the LLM: $q_{\text{LM}}(z_{\text{LM}}|m)$.
This allows us to frame the end-to-end process—from human experience to LLM representation—as a two-stage process of representation transformation, as illustrated in Figure 7.